OpenStudy (anonymous):

Find the slope of 4x - 5y=2

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

if y is the dependent variable.........

OpenStudy (anonymous):

ax+by+c=0 > slope is -a/b so, -4/-5 =4/5

OpenStudy (anonymous):

bring in form y=mx+c . Then the slope will become 4/5